in a math test,the mean is 45,and the standard deviation is 4.Assuming normality,what is the probability that a score picked at random will lie above score is 50



1
Expert's answer
2022-03-28T16:02:23-0400

"LP(X>50)=1-P(X<\frac{50-45}{4})=1-P(X<1.25)=1-0.89435=0.10565"